What mass of calcium chloride is required to make 500. mL of a 0.500M solution?
1
Expert's answer
2020-04-28T12:19:59-0400

V = 0.5 L M = 0.5 mol/L

M = n/V n = M*V

n = 0.5*0.5 = 0.25 moles

1 mol CaCl2 - 110.98 (g)

0.25 mol CaCl2 - x (g)

x = 27.745 (g)

Answer: 27.74 g.